OpenStudy (anonymous):

A triangle has an area of 8m^2+12m square units. Which expression represents the base of a triangle?

OpenStudy (anonymous):

oh really there is now way of knowing unless you have a picture. they want you to say \[8m^2+12m=4m(m+3)\]

OpenStudy (anonymous):

and area is \[\frac{1}{2}bh\]

OpenStudy (anonymous):

so \[\frac{1}{2}bh=4m(m+3)\] \[bh=8m(m+3)\]

OpenStudy (watchmath):

Is there any information what kind of triangle do you have? right triangle? equilateral?

OpenStudy (anonymous):

some math teacher made up this problem

OpenStudy (anonymous):

like that geometric series one where the first term was \[\sqrt{x}\] and the third term was \[4\sqrt{5}\] and you were supposed to deduce r as if it was possible